There are 35 students in Ms. Pool's first grade class. All of the students play either baseball or soccer, and some play both. If 11 of the students play baseball and 31 play soccer, how many play soccer but not baseball?
Answer by richard1234(7193)

Inclusion-exclusion
35 = 11 + 31 - (# students who play both)

# of students who play both = 7
# of students who play soccer but not baseball = 31 - 7 = 24